a) At which distance from a charge of q = +12nC should you place a charge of q = -3mC so that the potential energy is 5x10-3J?


b) Two identical charges q1 and q2 form a system with a potential energy of 2000J when the separation between them is of 0.5 m. Find the magnitude of each charge. 



1
Expert's answer
2021-10-28T15:21:30-0400

a)

"U=\frac{k|q_1q_2|}r,\implies r=\frac {k|q_1q_2|}U=64.8~m,"

b)

"U=\frac{kq^2}r,\implies q=\sqrt{\frac{Ur}k}=0.33~mC."